OpenStudy (anonymous):

A Sail Is In The Form Of A Right Triangle That Is Four Times As A High As It Is Wide. The Sail Is Made From 18 Square Meters Of Material. What Is The Height

OpenStudy (mayankdevnani):

area=1/2*base*height

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

So 1/2*18

OpenStudy (mayankdevnani):

18=1/2*b*h

OpenStudy (mayankdevnani):

h=2*b

OpenStudy (anonymous):

1/2*2

OpenStudy (mayankdevnani):

18=1/2*b*4b

OpenStudy (mayankdevnani):

got it

OpenStudy (anonymous):

No Im Still Stuck On The B

OpenStudy (anonymous):

Is It 3

OpenStudy (mayankdevnani):

breath is 4 times as base

OpenStudy (anonymous):

Okay So Will The B Equal To 3

OpenStudy (anonymous):

@hartnn Please Help

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

Okay So The Answer Is 3

OpenStudy (mayankdevnani):

yaa

OpenStudy (mayankdevnani):

@hartnn am i right

hartnn (hartnn):

yes

OpenStudy (mayankdevnani):

thnx...